OpenStudy (anonymous):

A woman drives 6 miles, accelerating uniformly from rest to 70 mph. How long does it take for her to reach 70 mph?

OpenStudy (anonymous):

Recall v = d/t v = velocity d = distance t = time Therefore: 70 = 6/t t = 6/70 hours

OpenStudy (anonymous):

lololol the above answer is so wrong, it does not factor in ACCELERATION

OpenStudy (anonymous):

lol you're right. Then that would make it a physics question. d = (vi+vf/2)t I may be wrong, I'm no good at physics.